OpenStudy (vshiroky):

@abb0t can you help please?

OpenStudy (vshiroky):

Suppose it is a hot summer day (40.0 degrees C = 104.0 degrees F), and you have left a can of hairspray outside so the gas within it is now at the same temperature of its surroundings. As part of your chemistry lab class, you have to build a safe and effective potato gun to share with the local elementary school, and you must be sure that the combustion chamber is well insulated and will not harm (burn) the shooter. You spray the warm hairspray into the 3.50 L chamber on your gun. You seal the cap, then ignite it, and the resulting gas expands to 13.0 L as the potato is shot out. What is the final temperature, in Celsius, inside the combustion chamber? What is this temperature in Fahrenheit?

OpenStudy (anonymous):

If the pressure remains constant then you can apply Charles' Law (although I'm unsure how this could be an isobaric process) which is \[\frac{ V _{1} }{ T _{1} } = \frac{ V_2 }{ T_2 } \] If there is some other information (such as the number of molecules of gas or amount of gas) which allows you to calculate the pressure inside the 2.50L chamber then you can apply the ideal gas law \[\frac{ p_1 V _{1} }{ T _{1} } = \frac{ p_2 V_2 }{ T_2 } \] \[T_2 = \frac{ p_1 V_1 }{ T_1 p_2 V_2 } \] Where \[p_1 = \frac{ nRT_1 }{ V_1 } \]

OpenStudy (vshiroky):

@iPwnBunnies @satellite73 I'm not sure how to tell what is what?

OpenStudy (vshiroky):

@Zale101fan

OpenStudy (vshiroky):

@thomaster

OpenStudy (vshiroky):

Please someone!!!

OpenStudy (vshiroky):

@anitasonia98

OpenStudy (vshiroky):

Please it's due in 30 minutes and I'm lost

OpenStudy (anonymous):

I would try: 3.50 L/313 K = 13.0 L/__ K (40 degrees Celsius in Kelvin is 313 K; you add 273 to your Celsius temp to get Kelvin. For any problem having to do with gas laws and such, your temperature must always be in Kelvin.)

OpenStudy (anonymous):

I hope that helps :)

OpenStudy (vshiroky):

Do I multiply straight across @anitasonia98

OpenStudy (anonymous):



OpenStudy (vshiroky):

I'm confused how to do that? I would have thought 3.5 times 13 divided by 313?

OpenStudy (anonymous):

You'd cross multiply, since you're dealing with proportions: 3.50x=(313)(12) :)

OpenStudy (anonymous):

If you solve for x, that should give you your answer :3

OpenStudy (vshiroky):

I got 1,162.57... I feel that is wrong?

OpenStudy (vshiroky):

did I do it wrong @anitasonia98

OpenStudy (anonymous):

I got 1073.14..

OpenStudy (vshiroky):

Isn't is 313(13) not 12?

OpenStudy (anonymous):

Oh, oops. Your answer is right xD

OpenStudy (vshiroky):

Ok.. what do I do next though?

OpenStudy (anonymous):

1162.57-273 Since your answer is in Kelvin, you have to convert it back into Celsius...So you subtract by 273 to get your Celsius temperature :3

OpenStudy (vshiroky):

Thank you and how do I convert to F

OpenStudy (anonymous):

Celsius temp x 9/5 -32

OpenStudy (vshiroky):

So C=889.57

OpenStudy (anonymous):

That seems right to me :)

OpenStudy (vshiroky):

I got 1,569.23 F

OpenStudy (anonymous):

Yup :)

OpenStudy (vshiroky):

Thank you!!!

OpenStudy (anonymous):

No prob :)